# Bootstrapping and autonomous tuning system for Si/SiGe multi-quantum dot devices

### v3

@@ III.1 First Leakage Test
When initiated, BATIS first checks the device for leakage to the fridge ground through the oxide or any other possible means in the instrumentation setup. Leakage can be caused by, e.g., defects arising during the fabrication or packaging processes.
The first leakage measurement happens before the electron accumulation, during the leakage test stage; see Fig. 2. The resistances of all the connections are measured with respect to each other and fridge ground, and they are compiled into a leakage matrix, as exemplified in the second and third row in Fig. 3(a). Leakage manifests as off-colored cells in the matrix, as shown in the right panel in the second row and in the third row of Fig. 3(a). The desired result is a high resistance between all the connections. Should leakage be detected, BATIS terminates and flags the device as unusable.
@@ III.2 Global Accumulation
For accumulation mode QD devices, no 2DEG is present at zero applied gate voltage. Applying a non-zero positive bias lowers the conduction band minimum below the Fermi level, accumulating a 2DEG in the quantum well near the Si/SixGe1-x interface. The goal of the global accumulation stage is to determine the characteristic turn-on voltage $V_{\mathcal{P}}$ where enough of the 2DEG is accumulated that a significant amount of current begins to flow between the ohmic contacts, effectively turning the device on; see Fig. 3(b). Since BATIS assumes no a priori knowledge about the device response, $V_{\mathcal{P}}$ is determined through an iterative approach, with $V_{\rm max}$ increased incrementally until BATIS is confident that $V_{\mathcal{P}}$ is observed in the measured window.
Idealized successful and failing measurements for this stage are visualized in the left and right panels in the middle row of Fig. 3(b), respectively. The turn-on process in real QD devices is not always straightforward, as even fully functional devices may fail to turn on correctly during initial testing. If the algorithm encounters a problem during global turn-on, it has a recovery pathway through optional illumination via a 780 $\mathrm{nm}$ light to adjust the microscopic configuration, magnitude, and sign of trapped charge, which can dramatically alter the device’s response to gate voltages [34]; see Fig. 2.
After illumination, the global accumulation stage is re-initiated. If the preference is not to use illumination, BATIS allows skipping this step and going straight to repeating the global turn-on stage. Whether or not illumination should be used is an option that is pre-defined in the configuration file.
Determining the turn-on voltages VP relies on fitting to a phenomenological function that mimics the expected behavior of the signal. Prior to fitting, the raw data is preprocessed to clean the signal. The preprocessing involves cropping the data to avoid non-linearities at the extreme limits of the data acquisition electronics, followed by smoothing the data to remove noise from the signal and normalization. We then parameterize the $I/V$ response of the device using
@@
where $V_{0}$ is the center of the step function and $\nicefrac{{1}}{{4\delta V}}$ is the slope at $V_{0}$. This function mimics the expected exponential behavior of the Fermi-Dirac distribution of electron energies with respect to gate voltages [40, 41, 42]. Similarly to the $I/V$ characteristics of the metal–oxide–semiconductor field-effect transistors (MOSFETs), an expected saturation will occur, leading to a roughly sigmoidal shape [43].
@@
${\rm V}_{\mathcal{P}}=V_{0}+8v\,\delta V,$ (2)
where v=0.5v=-0.5 is a fixed constant specifying how much the gate needs to be adjusted to turn-on in the units of $\delta V$  [^2] The parameter $v$ is defined as part of the configuration file and can be adjusted if needed.. The safe maximum allowed voltage applied to each gate, $V_{\rm max}$, is then updated to better match the saturation of the curve, such that it is less than or equal to the original. This voltage is later used as the starting point for the characterization of the screening and reservoir pinch-offs.
@@ III.3 Accumulated Leakage Test
With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test. This step proceeds analogously to the first leakage check, except that it does not include the ohmic connections. This is because this leakage check is taken when the device is accumulated, and the ohmics are intentionally shorted to each other. If this stage fails, it indicates that a gate is shorted to the 2DEG flowing beneath it, which is an unrecoverable fabrication defect. Should that occur, BATIS terminates.
@@ III.4 Screening and Reservoir Characterization
The derivative of the $I/V$ response signals for screening and reservoir gate pinch-offs are preprocessed in a similar fashion as described in Sec. III.2. The pinch-off, guaranteed isolation, and operating voltages are defined following Eq. (2), with v=0.5v=-0.5, v=1.0v=-1.0, and $v=0.5$, respectively. The pinch-off occurs at roughly the inflection of the signal VP, the guaranteed isolation, VI, happens left of the pinch-off, and the operating voltages, VO, occur to the right of the pinch-off, typically on the slope. The operating voltage for the central screening gate is handled separately as it has to satisfy a secondary constraint: all current channels must be formed without merging. By default, v=0.75v=-0.75. If at the subsequent stage a failure to form 1D current channels is observed twice, BATIS adjusts this parameter in increments defined as $v_{\tau}=(-1)^{p}\,0.02$, where $p=\{0,1\}$ is chosen at random.
Typical unsatisfactory operational modes for this stage include too much depletion or accumulation; see the second row, the right column of Fig. 3(c). If the channel below a gate is extraordinarily depleted, no voltage can be applied to the gate to accumulate it. On the other hand, if the channel is too accumulated, it is not possible to pinch it off. However, both types of issues would have been detected at the global accumulation stage, prompting the initiation of device illumination. Thus, at this stage, the constant current signal is caused by either an instrument failure or a broken gate. To verify whether this is the case, BATIS executes a repeated measurement. If the error is observed a second time, it indicates an issue with the gate and the algorithm terminates.
@@ III.5 Current Channel Accumulation
The 1D current channels are formed by selectively modifying the 2DEG in specific regions of the device. It involves depleting charges underneath the screening gates and accumulating them underneath the plunger and barrier gates (often collectively called finger gates) such that the current flows between the screening gates along the desired path, as seen in the top row of Fig. 3(d).
@@
Determining voltages where 1D channels are properly formed is typically done iteratively, by sweeping two screening gates against each other in a 2D measurement while holding the corresponding finger gates constant. The formation of the channel manifests as a triangular-sloped region (the so-called triangle plots[45]. If the finger gates are not sufficiently accumulated, this region will fail to form. Multiple 2D scans with incremented finger gate voltages for each pair of screening gates—a three-dimensional dataset—are often necessary to observe the triangle region across all required channels.
@@
Figure 4: Interplay between screening and finger gates during the single channel formation. (a) The current continuity simulation and (b) an example of an experimentally acquired $I/V$ response landscape in the screening gate vs. finger gates space. (c) A Schrödinger-Poisson simulated example showing three different operation modes, with each current channel programmed to be at a different operation mode: I. A proper 1D channel. II. An improper 2D channel under the screening and finger gates. III. An improper 2D channel under the screening gate. (d) A Schrödinger-Poisson simulated example of the proper channel formation operation point for the device. Details about simulations are in the Supplemental Material [37].
@@
BATIS employs an alternative approach for initializing multiple 1D channels simultaneously. Rather than sweeping screening gates against each other, BATIS performs a 2D measurement of a screening gate against all the finger gates located between that gate and the opposing screening gate used to define the 1D channel (typically the central screening gate). Thus, this method efficiently captures the behavior of all channels in a single 2D dataset.
@@
A simulation of current as a function of gate voltages is shown in Fig. 4(a), and in Fig. 4(b) a corresponding example of experimental data is shown. Notably, there are four unique regions in the landscape in Fig. 4(a): a no-current region in the bottom left, a finger-gate-dominated region in the top left (indicated with I), a screening-gate-dominated region in the bottom right (indicated with III), and a region of combined effect from both the fingers and screening gates (indicated with II). These regions all correspond to different operating modes of the device, with region I showing the proper accumulation of a 1D channel that is the required precursor to proper bootstrapping of the device; region II, in contrast, corresponds to an improper 2D channel formed under the screening and finger gates; and region III corresponds to an undesired 2D channel under the screening gate. All three of these regions are indicated in the simulation results shown in Fig. 4(c). If the finger-gate-dominated region does not form, as shown in the right panel in the second row of Fig. 3(d), BATIS initiates illumination with finger gates voltages set more negative than the most recent illumination.
@@
When the finger-gate-dominated region forms properly, BATIS automatically selects an operating point in region I, near the boundary between the current and the no-current regions. Fig. 4(d) shows a Schrodinger-Poisson simulated example with properly set operating points for all current channels. The 1D channel formation technique implemented in BATIS requires a single measurement for each pair of screening/finger gates, which makes it much more efficient than the standard approaches.
@@ III.6 Finger Gate Characterization
Once isolated current channels are formed, finger gate pinch-offs, VP, can be measured. Naively, one might hope that, as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend like that shown in the middle row of Fig. 3(e). However, because finger gates are so small, trapped charge, e.g., in the oxide layer, often causes Coulomb blockade effects [26] to be visible, so that real devices display behavior like that shown in the third row of Fig. 3(e). In this case, as the gate voltage is made less positive, the curve exhibits characteristic Coulomb blockade oscillations typical of QD devices. To be robust against this non-ideality, BATIS uses thresholding and slope detection during analysis to determine the pinch-off voltage on a point-by-point basis, going in the direction of increasing voltage from the end of the sweep (the reverse of how the signal was measured) to avoid small deviations of background noise, while capturing the first significant turn-on behavior.
@@ IV BATIS: Experimental Demonstration
We demonstrate BATIS at 1.3 $\mathrm{K}$ on a Si/SixGe1-x quadruple QD device with 24 gates, nominally identical to the one shown in Fig. 1(a), which has two charge sensing channels in parallel with a channel to host the four qubit QDs. The first gate layer contains four screening gates, Si for $i=1,\cdots,4$, shown in Fig. 1(a) in yellow. The next layer consists of five reservoirs, Ri for $i=1,\cdots,5$, and six plunger gates, Pi for $i=1,\cdots,6$. The reservoirs are shown in blue in Fig. 1(a), while the gates highlighted in red are the plunger gates. The final layer consists of nine barrier gates, Bi for $i=1,\cdots,9$, shown in Fig. 1(a) in gray.
The goal of BATIS is to form three current channels and determine proper pinch-off gate voltages based on automated measurements and the gate layout information BATIS learns from the configuration file. This proper operation is indicated with black arrows in Fig. 1(b), bringing a grounded but not-characterized device to a fully operational mode. In what follows, we explain how BATIS navigates the execution of the quad-QD device bootstrapping, testing, and characterization.
First Leakage Test. The sample mount holding the quad-QD device used here contains $40$ connections, of which $29$ are used for instrument channels to control individual gates and ohmics, and an additional $11$ wired channels are unused for this device. Thus, the size of the leakage matrix is $40\times 40$ to capture all possible leakage states. The leakage matrix is symmetric, and thus only the lower triangular part needs to be completed.
BATIS first checks the $40$ diagonal elements of the leakage matrix. If the resistance of any connection drops below 25 $\mathrm{M\SIUnitSymbolOhm}$, BATIS declares that the device leaks. Moreover, if the threshold test fails on a diagonal component, BATIS automatically fills out the rest of the column for diagnostic purposes. In the worst case, a total of $820$ resistance measurements are necessary to complete the leakage matrix.
@@
Table 1: The summary of the four global accumulation runs. The first column indicates the channel, and the derived voltages are indicated in the second column. The VP indicates the detected turn-on, while Vmax is the maximum safe accumulation voltage for a given channel. BATIS chose to initiate illumination after the first three runs. The fourth run was satisfactory, and the algorithm proceeded to the next stage.
Voltages ($\mathrm{mV}$) Run 1 Run 2 Run 3 Run 4
I1 VP $28.0$ $26.1$ $24.7$ $296.0$
Vmax $81.7$ $78.6$ $75.5$ $349.4$
I2 VP $28.2$ $20.8$ $23.5$ $308.7$
Vmax $53.6$ $47.3$ $47.3$ $339.2$
I3 VP $7.4$ $9.5$ $2.6$ $317.1$
Vmax $53.6$ $53.6$ $50.4$ $349.4$
Global Accumulation. Next, BATIS seeks to determine gate voltages that establish global accumulation. This step generates current flow across the device without regard to which channel or channels are connected to each other. In our demonstration, the algorithm executed three rounds of illumination before the observed global turn-on values were determined to be satisfactory, occurring in the desired range of $200$ $\mathrm{mV}$ to $400$ $\mathrm{mV}$. By design, all gates are grounded during the first two illuminations. The first illumination is intended to completely reset the device, while the second is to verify that all instruments work as intended. Because BATIS still did not find the gates in the desired range for the third illumination, voltages were applied to all screening and reservoir gates. The voltages required to achieve turn-ons in the required range are determined by BATIS following the protocol in Ref. [34]. In this case, during illumination, BATIS sets gates associated with channels I1 and I2 to $280$ $\mathrm{mV}$, and gates associated with channel I3 to $300$ $\mathrm{mV}$. The central screening gate, S2, was left grounded (i.e., set to 0 $\mathrm{mV}$) during illumination.

### v4

@@ III.1 First leakage test
When initiated, BATIS first checks the device for leakage to the fridge ground through the oxide or any other possible means in the instrumentation setup. Leakage can be caused, for example, by defects arising during fabrication or packaging.
The first leakage measurement happens before the electron accumulation, during the leakage test stage; see Fig. 2. The resistances of all the connections are measured with respect to each other and the fridge ground, and they are compiled into a leakage matrix, as exemplified in the second and third rows in Fig. 3(a). Leakage manifests itself as off-colored cells in the matrix, as shown in the right panel in the second row and in the third row in Fig. 3(a). The desired result is a high resistance between all the connections. Should leakage be detected, BATIS terminates and flags the device as unusable.
@@ III.2 Global accumulation
For accumulation-mode QD devices, no 2DEG is present at zero applied gate voltage. Application of a nonzero positive bias lowers the conduction band minimum below the Fermi level, resulting in accumulation of a 2DEG in the quantum well near the Si/SixGe1-x interface. The goal of the global accumulation stage is to determine the characteristic turn-on voltage $V_{\mathcal{P}}$ where enough of the 2DEG is accumulated that a significant amount of current begins to flow between the Ohmic contacts, effectively turning the device on; see Fig. 3(b). Since BATIS assumes no a priori knowledge of the device response, $V_{\mathcal{P}}$ is determined through an iterative approach, with $V_{\rm max}$ increased incrementally until BATIS is confident that $V_{\mathcal{P}}$ is observed in the measured window.
Idealized successful and failing measurements for this stage are visualized in the left and right panels, respectively, in the second row in Fig. 3(b). The turn-on process in real QD devices is not always straightforward, as even fully functional devices may fail to turn on correctly during initial testing. If the algorithm encounters a problem during global turn-on, it has a recovery pathway through optional illumination via 780 $\mathrm{nm}$ light to adjust the microscopic configuration, magnitude, and sign of trapped charge, which can dramatically alter the device’s response to gate voltages [34]; see Fig. 2.
After illumination, the global accumulation stage is reinitiated. If the preference is not to use illumination, BATIS allows skipping of this step and proceeding directly to repeating the global turn-on stage. Whether or not illumination should be used is an option that is predefined in the configuration file.
Determination of the turn-on voltages $V_{\mathcal{P}}$ relies on fitting to a phenomenological function that mimics the expected behavior of the signal. Before fitting, the raw data are preprocessed to clean the signal. The preprocessing involves cropping the data to avoid nonlinearities at the extremes of the data-acquisition electronics, followed by smoothing to remove noise from the signal and normalization. We then parameterize the II-$V$ response of the device using
@@
where $V_{0}$ is the center of the step function and $\nicefrac{{1}}{{4\delta V}}$ is the slope at $V_{0}$. This function mimics the expected exponential behavior of the Fermi-Dirac distribution of electron energies with respect to gate voltages [40, 41, 42]. Similarly to the II-$V$ characteristics of metal–oxide–semiconductor field-effect transistors, an expected saturation will occur, leading to a roughly sigmoidal shape [43].
@@
$V_{\mathcal{P}}=V_{0}+8v\,\delta V,$ (2)
where v=0.5v=-0.5 is a fixed constant specifying how much the gate needs to be adjusted to turn on in the unit of $\delta V$  [^2] The parameter $v$ is defined as part of the configuration file and can be adjusted if needed.. The safe maximum allowed voltage applied to each gate, Vmax, is then updated to better match the saturation of the curve, such that it is less than or equal to the original. This voltage is later used as the starting point for the characterization of the screening and reservoir pinch-offs.
@@ III.3 Accumulated leakage test
With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test. This step proceeds analogously to the first leakage check, except that it excludes the Ohmic connections. This is because this leakage check is performed when the 2DEG in the device is accumulated, with the Ohmic connections intentionally shorted to each other. If this stage fails, it indicates that a gate is shorted to the 2DEG flowing beneath it, which is an unrecoverable fabrication defect. Should that occur, BATIS terminates.
@@ III.4 Screening and reservoir characterization
The derivative of the II-$V$ response signals for screening and reservoir gate pinch-offs are preprocessed in a similar fashion as described in Sec. III.2. The pinch-off, guaranteed isolation, and operating voltages are defined following Eq. (2), with v=0.5v=-0.5, v=1.0v=-1.0, and $v=0.5$, respectively. The pinch-off occurs at roughly the inflection of the signal $V_{\mathcal{P}}$, the guaranteed isolation VI happens left of the pinch-off, and the operating voltages VO occur to the right of the pinch-off, typically on the slope. The operating voltage for the central screening gate is handled separately, as it has to satisfy a secondary constraint: all current channels must be formed without merging. By default, v=0.75v=-0.75. If at the subsequent stage a failure to form 1D current channels is observed twice, BATIS adjusts this parameter in increments defined as $v_{\tau}=(-1)^{p}\,0.02$, where $p=\{0,1\}$ is chosen at random.
Typical unsatisfactory operational modes for this stage include too much depletion or accumulation; see the right panel in the second row in Fig. 3(c). If the 2DEG along the channel below a gate is extraordinarily depleted, no voltage can be applied to the gate to accumulate it. On the other hand, if the channel is too accumulated, it cannot be pinched off. However, both types of issue would have been detected at the global accumulation stage, prompting the initiation of device illumination. Thus, at this stage, the constant current signal is caused by either an instrument failure or a broken gate. To verify this, BATIS conducts a repeated measurement. If the error is observed a second time, this indicates an issue with the gate, and the algorithm terminates.
@@ III.5 Current channel accumulation
The 1D current channels are formed by selectively modifying the 2DEG in specific regions of the device. This involves depleting charges underneath the screening gates and accumulating them underneath the plunger and barrier gates (often collectively called finger gates) such that the current flows between the screening gates along the desired path, as seen in the first row of Fig. 3(d).
@@
Determination of voltages at which 1D channels are properly formed is typically done iteratively, by sweeping two screening gates against each other in a 2D measurement while holding the corresponding finger gates constant. The formation of the channel manifests itself as a triangular-sloped region (the so-called triangle plots[45]. If the 2DEG under the finger gates along the channel is not sufficiently accumulated, this region will fail to form. Multiple 2D scans with incrementally increased finger-gate voltages for each pair of screening gates—a three-dimensional dataset—are often necessary to observe the triangle region across all required channels.
@@
Figure 4: Interplay between screening and finger gates during the single channel formation. (a) Current continuity simulation and (b) an example of an experimentally acquired II-$V$ response landscape in the screening gate vs finger gate space. (c) Schrödinger-Poisson simulated example showing three different operation modes, with each current channel programmed to be at a different operation mode: I, a proper 1D channel; II, an improper 2D channel under the screening and finger gates; III, an improper 2D channel under the screening gate. (d) Schrödinger-Poisson simulated example of the proper channel formation operation point for the device. Details about simulations are given in Supplemental Material [37].
@@
BATIS uses an alternative approach to simultaneously initialize multiple 1D channels. Rather than sweeping screening gates against each other, BATIS performs a 2D measurement of a screening gate against all the finger gates located between that gate and the opposing screening gate used to define the 1D channel (typically the central screening gate). Thus, this method efficiently captures the behavior of all channels in a single 2D dataset.
@@
A simulation of current as a function of gate voltages is shown in Fig. 4(a), and in Fig. 4(b) a corresponding example of experimental data is shown. Notably, there are four unique regions in the landscape in Fig. 4(a): a no-current region at the bottom left, a finger-gate-dominated region at the top left (indicated by “I”), a screening-gate-dominated region at the bottom right (indicated with “III”), and a region of combined effect from both the finger gates and the screening gates (indicated by “II”). These regions all correspond to different operating modes of the device, with region I showing the proper accumulation of a 1D channel that is the required as a precursor to proper bootstrapping of the device; region II, in contrast, corresponds to an improper 2D channel formed under the screening and finger gates; and region III corresponds to an undesired 2D channel under the screening gate. All three of these regions are indicated in the simulation results shown in Fig. 4(c). If the finger-gate-dominated region does not form, as shown in the right panel in the second row in Fig. 3(d), BATIS initiates illumination with finger gate voltages set more negative than for the most recent illumination.
@@
When the finger-gate-dominated region forms properly, BATIS automatically selects an operating point in region I, near the boundary between the current region and the no-current region. Figure 4(d) shows a Schrödinger-Poisson simulated example with properly set operating points for all current channels. The 1D-channel-formation technique implemented in BATIS requires a single measurement per pair of screening and finger gates, making it much more efficient than standard approaches.
@@ III.6 Finger Gate Characterization
Once isolated current channels are formed, finger gate pinch-offs $V_{\mathcal{P}}$ can be measured. Naively, one might hope that as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend like that shown in the second row in Fig. 3(e). However, because finger gates are so small, trapped charge (e.g., in the oxide layer) often causes Coulomb blockade effects [26] to be visible, so real devices display behavior like that shown in the third row in Fig. 3(e). In this case, as the gate voltage is made less positive, the curve exhibits characteristic Coulomb blockade oscillations typical of QD devices. To be robust against this nonideality, BATIS uses thresholding and slope detection during analysis to determine the pinch-off voltage on a point-by-point basis, going in the direction of increasing voltage from the end of the sweep (the reverse of how the signal was measured) to avoid small deviations of background noise, while capturing the first significant turn-on behavior.
@@ IV BATIS: Experimental Demonstration
We demonstrate BATIS at 1.3 $\mathrm{K}$ on a Si/SixGe1-x quadruple QD device with 24 gates, nominally identical to the one shown in Fig. 1(a), which has two charge-sensing channels in parallel with a channel to host the four qubit QDs. The first gate layer contains four screening gates (S1-S4) shown in Fig. 1(a) in yellow. The next layer consists of five reservoirs (R1-R5) and six plunger gates (P1-P6). The reservoirs are shown in blue in Fig. 1(a), while the gates highlighted in red are the plunger gates. The final layer consists of nine barrier gates (B1-B9) shown in Fig. 1(a) in gray.
The goal of BATIS is to form three current channels and determine appropriate pinch-off gate voltages on the basis of automated measurements and the gate layout information BATIS learns from the configuration file. This proper operation is indicated by black arrows in Fig. 1(b), bringing a grounded but not-characterized device to a fully operational mode. In what follows, we explain how BATIS navigates the execution of the bootstrapping, testing, and characterization of the quadruple QD device.
IV.1 First leakage test
The sample mount holding the quadruple QD device used here contains $40$ connections, of which $29$ are used for instrument channels to control individual gates and Ohmic contacts, and an additional $11$ wired channels are unused for this device. Thus, the leakage matrix is $40\times 40$ to capture all possible leakage states. The leakage matrix is symmetric, and thus only the lower triangular part needs to be completed.
BATIS first checks the $40$ diagonal elements of the leakage matrix. If the resistance of any connection drops below 25 $\mathrm{M\SIUnitSymbolOhm}$, BATIS declares that the device leaks. Moreover, if a diagonal component fails the threshold test, BATIS automatically fills out the rest of the column for diagnostic purposes. In the worst case, a total of $820$ resistance measurements are necessary to complete the leakage matrix.
@@ IV.1 First leakage test
IV.2 Global accumulation
[…]
Following the gate-biased illumination, the fourth iteration of the global turn-on measurement was successful. A summary of all four runs is presented in Table 1. The resulting configuration of the QD device is shown in Fig. 5(a), where the pinch-off voltages on all the screening gates, excluding S2, and all the reservoir gates are colored on the basis of the learned gate voltage.
Table 1: Summary of the four global accumulation runs. $V_{\mathcal{P}}$ represents the detected turn-on voltage, while $V_{\mathrm{max}}$ is the maximum safe accumulation voltage for a given channel. BATIS chose to initiate illumination after the first three runs. The fourth run was satisfactory, and the algorithm proceeded to the next stage.
Voltages Run 1 Run 2 Run 3 Run 4
I1 $V_{\mathcal{P}}$ ($\mathrm{mV}$) $28.0$ $26.1$ $24.7$ $296.0$
$V_{\mathrm{max}}$ ($\mathrm{mV}$) $81.7$ $78.6$ $75.5$ $349.4$
I2 $V_{\mathcal{P}}$ ($\mathrm{mV}$) $28.2$ $20.8$ $23.5$ $308.7$
$V_{\mathrm{max}}$ ($\mathrm{mV}$) $53.6$ $47.3$ $47.3$ $339.2$
I3 $V_{\mathcal{P}}$ ($\mathrm{mV}$) $7.4$ $9.5$ $2.6$ $317.1$
$V_{\mathrm{max}}$ ($\mathrm{mV}$) $53.6$ $53.6$ $50.4$ $349.4$
IV.3 Accumulated Leakage test
With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test. A second leakage matrix is measured, following the same procedure as during the first leakage test. Like previously, BATIS finds that all resistances exceed 25 $\mathrm{M\SIUnitSymbolOhm}$. Since none of the connections leaked, only $35$ measurements following the matrix diagonal were taken.